OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN! explain why (4, 1) is not a solution to the equation y = 3x + 1

OpenStudy (swissgirl):

Well lets plug it in When x=4 does y=1??? The only way to find out is to plug it into the equation y=3(4)+1 =12+1 =13 So when x=4 then y=13 So therefore we can conclude that this line doesnt pass through the point (4,1)

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} (&4,&1)\\ &{\color{brown}{ x}}&{\color{blue}{ y}}\\ \end{array}\implies {\color{blue}{ y}}=3{\color{brown}{ x}}+1\)

OpenStudy (anonymous):

Thank you both! :)